# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

### v2

@@ Appendix E Comprehensive Comparison with Other Methods
In order to comprehensively evaluate the performance of our proposed solution, we compare Mujica-MyGO with a series of RAG systems. Apart from direct inference with pretrained LLM [78], we consider vanilla RAG [29], RAFT [87], RAFe [45], HippoRAG [17], Iter-RetGen [57], IRCoT [68], IterDRAG [83], as well as compatitative concurrent solutions Search-o1 [34], RAG-Gym [29] and ReSearch [10]. Considering the differences between the backbone LLMs and retrieval methods employed by each RAG system, it is inappropriate to compare the performance of different solutions solely based on their evaluation metrics.
Table 4 presents the experimental results on three representative multi-hop question answering datasets. We have the following observations:
@@
Firstly, Mujica-MyGO is an effective and efficient multi-hop QA solution. Despite leveraging a 7B Qwen2.5 LLM as its backbone, Mujica-MyGO achieves state-of-the-art performance on the MuSiQue and 2Wiki datasets and delivers competitive results on HotpotQA. One may notice that some baseline methods outperform Mujica-MyGO on the HotpotQA dataset. However, these approaches typically rely on larger-scale backbone LLMs and/or more sophisticated retrieval methods. In general, larger backbone models (e.g., GPT and QwQ) tend to yield better results. In addition, incorporating advanced retrievers (e.g., HopRetriever [33], Dragon-plus [37]) or expanding the knowledge corpus via search engines (e.g., Google, Bing) can further enhance performance. Nevertheless, it is important to emphasize that improving retrieval is not the focus of this paper. In order to evaluate the effectiveness of the proposed method, we opt to adopt the official corpus released with the HotpotQA dataset [80].
Secondly, Mujica shows to be an exceptional multi-hop QA pipeline. The results of the Mujica w/ warm up variant demonstrate that even without relying on expensive reinforcement learning (RL), a small amount of fine-tuning data is sufficient for Mujica to outperform RAG baselines based on question decomposition and iterative retrieval. We attribute the advantages of our method to two key factors. On the one hand, the use of independent workers to process retrieved contexts effectively prevents irrelevant information from interfering with the planner’s reasoning process. On the other hand, formulating subquestions into a DAG fully accounts for the interdependencies between subquestions, ensuring reasoning coherence throughout the pipeline. However, every approach has its trade-offs. As a complex question-answering pipeline, Mujica imposes substantial demands on the LLM’s instruction-following and reasoning capabilities, making it difficult for small-scale pre-trained LLMs to achieve desirable results in few-shot settings.
Table 5: Comparison of Proximity Optimization-based Multi-hop QA Solutions.
Pipeline Post-training Method HotpotQA MuSiQue 2Wiki
@@
Lastly, the experimental results reconfirm the effectiveness of MyGO. Compared with ReSearch (using GRPO), Search-R1 and RAG-Gym (using PPO), Mujica-MyGO achieves comparable or even better performance across three datasets with the commonly adopted bge-large-en-v1.5 dense retriever (see Table 5). This demonstrates that MyGO simplifies GRPO and PPO without compromising performance on multi-hop QA tasks. Moreover, MyGO does not require additional resources to hold a reference model or to compute advantages, striking a great balance between efficiency and performance.
Detailed comparisons between MyGO and other preference optimization algorithms (e.g., DPO, PPO, GRPO, etc.) across various tasks would be very interesting, however, this lies beyond the scope of this paper. We shall emphrasize that the focus of this paper is the multi-hop QA task. In our future works, we plan to extend our analysis to different reasoning tasks, including but not limited to mathematical problems, coding, and long-context reasoning.
Appendix F Further Study: Offline MyGO?
@@
Figure 4: F1 score on 2Wiki-KG training
Figure 5: F1 score on 2Wiki-Text training
Table 6: Mujica+MyGO w/ Qwen Online v.s. Offline on 2Wiki-KG and 2Wiki-Text Environments.
@@
We note that our goal is not to compare online and offline RL, as this has been discussed in many prior works [65]. Instead, we just aim to present an alternative approach for implementing MyGO.
@@
Appendix G Extended Related Works
@@
Retrieval-based methods dominate the landscape of textual MHQA since supporting contexts from external corpora can effectively compensate for deficiencies in the inherent knowledge of language models and mitigate hallucinations. Traditional approaches primarily utilized sentence transformers [8] for dense retrieval or BM25 [52] for keyword matching. Recently, extensive research has focused on optimizing the QA pipeline, particularly through the decomposition of complex questions [79, 81] and the iterative retrieval of relevant evidences in multiple steps [5, 22, 57, 68]. Most notably, ReAct [81] integrates retrieval actions into the reasoning trajectory exploration process, enabling dynamic interaction with external environments. IM-RAG [79] optimizes multi-round retrieval and reasoning via reinforced learning. Search-o1 [34] introduces an agentic RAG framework, empowering the language model to decide when to retrieve external knowledge and to refine retrieved contexts, achieving state-of-the-art performance.
@@
Early approaches [27, 49, 71] formulates KBQA as a semantic parsing task, focusing on extracting entities and relations from natural language questions and translating them into SPARQL statements capable of executing logical queries over KGs [59, 2]. While these methods provide accurate and interpretable answers, they encountered significant limitations since SPARQL statements generated by language models often suffer from syntactic errors, including incorrect entity and relation labels as well as unexpected grammatical mistakes [82, 42, 30]. Moreover, the incomplete nature of KGs [32] can also lead to query failures, preventing these approaches from outputting correct answers.
@@
Recently, researchers have shifted towards treating KBQA as an information retrieval (IR) task, leveraging RAG and the advanced semantic understanding capabilities of LLMs to generate answers based on factual knowledge stored in KGs [73]. LLM-based KBQA methods can be generally categorized into two types: iterative exploration and subgraph reasoning. Iterative exploration methods [63, 44, 86, 88, 9, 61, 77, 15, 64, 85, 30] adopt a beam search-like strategy to retrieve the most relevant relations and neighboring entities, followed by pruning operations guided by LLMs. These methods exhibit strong traceability, enabling responsible and interpretable reasoning, but often come with the drawback of high computational complexity [36]. Moreover, the absence of ground truth reasoning paths and intermediate nodes in most KBQA datasets limits the feasibility of applying SFT [42, 47]. As a result, their performance is inevitably constrained by the inherent reasoning capabilities of LLMs. In contrast, subgraph reasoning methods [42, 47, 31, 76, 43] retrieve all candidate reasoning paths related to the question from the local neighborhood of the topic entities and guide the LLM to generate answers based on these paths in a single step. This approach enable direct improvement of the LLM’s in-context learning capabilities through SFT. Nevertheless, its effectiveness is often constrained by the quality of retrieved context, which typically cover only a subset of the knowledge required to address complex questions, omitting other critical information [9]. Furthermore, due to the incompleteness of KGs, the existence of relevant reasoning paths within the subgraph is not guaranteed. In such scenarios, LLM may hallucinate and produce incorrect answers.
Simplifying Reinforcement Learning with LLMs
@@
Recent research has focused on simplifying RL algorithms for LLMs. For instance, GRPO [58] simplifies PPO by replacing the value model with standard deviation. Similarly, ReMax [35] adapts VPO by using the logits of greedy decoding as a baseline for GAE. Other VPO modifications include RLOO [4], REINFORCE++[20], and GPG [12]. Another line of research replaces VPO with MLE through advanced sampling techniques. Within this category, RAFT [87] selects the highest-reward trajectory within a group, while RAFT++ [75] enhances this by incorporating rejection sampling. Both RAFT and RAFT++ can be viewed as specific instances of STaR [84]. MyGO aligns with this latter research direction. It selects trajectories using either a fixed or progressively improving threshold, a property ensuring that sampling approaches the optimal policy. We further provide a mathematical theory to explain the efficacy of this streamlined method. If we continue to simplify MyGO, it will converge to STaR [84]. However, our primary motivation stems from the goal of simplifying RL algorithms specifically for LLM agent training.
Appendix H Working Examples
@@
Table 7: Prompt template of Planner for text-based multi-hop question answering
@@
Before answering, outline your plan and reasoning process within ‘<think>‘ tags. This includes understanding the question, identifying needed information, and planning your search queries.
@@
Please put your search query in ‘<search>‘ tags, e.g., ‘<search> Who is the mother of Xawery Żuławski? </search>‘
Every ‘<search>‘ tag only accepts one search query.
But you can use multiple ‘<search>‘ tags if needed.
@@
For example, if you directly search ‘Who is the mother of the director of film Polish-Russian War‘, the search engine cannot find the information directly.
@@
Table 8: Prompt Template of Worker for Text-based Multi-hop Question Answering
@@
$[0]$ Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University. She made her acting debut in 1984, and has since become active in television dramas, most notably as the arrogant and ambitious Lady Choi in the 2003 period drama "Dae Jang Geum" (or "Jewel in the Palace"), which was a hit not only in Korea but throughout Asia.
$[1]$ Shin Kyeong-nim was born on April 6, 1936 in North Chungcheong Province, South Korea. When he was young, Shin Kyong-rim frequented the people of Koreaś rural villages and collected the traditional songs they sang. Much of his poetry represents a modernization of things he heard then Shin Kyeong-nim graduated in English Literature from Dongguk University, from which time he strove to become a creative writer. In 1955 and 1956, he made his formal literary debut with the publication of poems "Day Moon" (Natdal), "Reeds" (Galdae) and "Statue of Stone" (Seoksang). He taught elementary school in his hometown for a period of time, before moving to Seoul to work as an editor for Hyundae munhak and Donghwa Publishers. But his strong desire to create poetry continued.
@@
$[3]$ Han Seung-yeon was born on July 24, 1988, in Seoul, South Korea. She made her acting debut as a child actress in a bit part in "Dear Ends" (1993), "Summer Showers" (1995) and "Star in My Heart" (1997). She later left South Korea to study at Tenafly High School in New Jersey, United States. However, she withdrew from high school mid-course in order to pursue a singing career. After returning to South Korea, she debuted as a member of girl group Kara on March 29, 2007. During her time with the group, she passed a high school qualification exam, the College Scholastic Ability Test, and was accepted by Kyung Hee University, majoring in theater and film.
@@
Passage [0] clearly states that "Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University."
@@
Table 9: Prompt Template of Planner for Knowledge-based Question Answering
@@
Please think thoroughly and label the final answer with <answer> *** </answer>.‘
@@
Search([0], "When did Blind Shaft come out?")
Search([1], "When did The Mask Of Fu Manchu come out?")
@@
Table 10: Prompt Template of Worker for Knowledge-based Question Answering
@@
First, analyze the question to identify the key entity and requested information (e.g., "mother," "birth date").

### v3

@@ 4.2.2. Effectiveness of Mujica
In order to comprehensively evaluate the performance of our proposed solution, we compare Mujica-MyGO with a series of RAG systems. Apart from direct inference with pretrained LLM Qwen2 and vanilla RAG RAG, we consider 7 iterative RAG solutions: RAFT zhang2024raftadaptinglanguagemodel, RAFe RaFe, HippoRAG gutiérrez2024hipporag, Iter-RetGen Iter-RetRAG, IRCoT IRCoT, IterDRAG IterDRAG, RAG-star RAG-Star and 4 RAG agents: Search-o1 Search-o1, RAG-Gym RAG and ReSearch ReSearch. Considering the differences in backbone LLMs and retrieval methods employed by each RAG system, it is inappropriate to compare the performance of different RAG systems solely based on their evaluation metrics.
@@
Table 3 presents the experimental results on three representative multi-hop question answering datasets, which shows that Mujica-MyGO is an effective and efficient multi-hop QA solution. Despite leveraging a 7B Qwen2.5 LLM as its backbone, Mujica-MyGO achieves state-of-the-art performance on theMuSiQue and 2Wiki datasets and delivers competitive results on HotpotQA. One may notice that some baseline methods outperform Mujica-MyGO on the HotpotQA dataset. However, these approaches typically rely on larger-scale backbone LLMs and/or more sophisticated retrieval methods. In general, larger backbone models (e.g., GPT and QwQ) tend to yield better results. In addition, incorporating advanced retrievers (e.g., HopRetriever HopRetriever, Dragon-plus DRAGON) or expanding the knowledge corpus via search engines (e.g., Google, Bing) can further enhance performance. Nevertheless, it is important to emphasize that improving retrieval is not the focus of this paper. In order to evaluate the effectiveness of the proposed method, we opt to adopt the official corpus released with the HotpotQA dataset yang2018hotpotqa. Across model variants, we find that even without costly reinforcement learning, a small amount of SFT data is sufficient for Mujica to surpass baseline methods based on question decomposition and iterative retrieval, the method to consistently surpass, demonstrating its robustness. However, every approach has its trade-offs. As a complex question-answering pipeline, Mujica imposes substantial demands on the LLM’s instruction-following and reasoning capabilities, making it difficult for small-scale pre-trained LLMs to achieve desirable results in few-shot settings.
4.3. Additional Studies
4.3.1. Effectiveness in Out-Of-Domain Settings
[…]
Type Model Hotpot $\to$ MuSiQue 2Wiki $\to$ QALD
EM F1 EM F1
Few-Shot
Qwen
2.5-7B
5.34 17.81 27.92 37.11
MyGO 26.11 35.92 39.85 49.73
4.3.2. MyGO v.s. Other Proximity Optimization Approaches
Compared with ReSearch (using GRPO), Search-R1 and RAG-Gym (using PPO), Mujica-MyGO achieves comparable or even better performance across three datasets with the commonly adopted bge-large-en-v1.5 dense retriever (see Table 6). This demonstrates that MyGO simplifies GRPO and PPO without compromising performance on multi-hop QA tasks. Moreover, MyGO does not require additional resources to hold a reference model or to compute advantages, striking a great balance between efficiency and performance.
Table 6. Comparison of Proximity Optimization-based Multi-hop QA Solutions.
Pipeline P.O. Method HotpotQA MuSiQue 2Wiki
@@ 4.3.1. Effectiveness in Out-Of-Domain Settings
4.3.3. Runtime Comparison - MyGO v.s. GRPO
We further present a runtime comparison between MyGO and GRPO. As shown in Table 7, Mujica trained with MyGO achieves performance comparable to that obtained with GRPO, while requiring approximately 74% less training time and about 94% fewer training trajectories.
Table 7. Comparison to GRPO. # of t.d.: means the number of trajectories for the planner and the worker; t. time: denotes the total training time.
Algo. F1 EM # of t.d. t. time
GRPO 86.06 80.15 137,941 $\sim$135min
MyGO 85.70 80.30 7,343 $\sim$35min
5. Related Work
MHQA with textual corpus.
Knowledge-based question answering.
6. Conclusion
In this work, we introduce Mujica, an interactive multi-agent framework for RAG systems, alongside MyGO, a novel reinforcement learning method designed to simplify its optimization. Mujica employs a planner-worker architecture to systematically decompose and resolve complex questions, while MyGO enhances training efficiency through an intelligent sampling strategy. Our experiments demonstrate that the proposed framework not only achieves significant performance improvement across multiple QA datasets but also exhibits strong inductive reasoning. These results underscore its scalability and potential to generalize effectively in real-world scenarios.
Appendix A Problem Definition
Text-based Multi-hop Question Answering
Knowledge-based Question Answering
Appendix B Mathematical Proofs
[…]
Proof.
Recall that $\pi^{*}(\tau)=\frac{\exp(r(\tau)/\alpha)}{Z(\alpha)}$ and $Z(\alpha)=\int\exp(r(\tau)/\alpha)d\tau$, we have the following statement for $\pi^{>K}$:
$\pi^{>K}(\tau)=\frac{\exp(r(\tau)/\alpha)}{Z^{>K}(\alpha)}\;,\quad\text{where }Z^{>K}(\alpha)=\int_{\mathcal{T}}\mathds{1}(r>K)\exp(r(\tau^{\prime})/\alpha)d\tau^{\prime}\;,$ (7)
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})=\log\frac{Z}{Z^{>K}}\;.$ (8)
Notably,
$Z^{>K}=\int_{r>K}\exp(r/\alpha)dr=(M-K)\exp(r^{\prime}/\alpha)\;,$ (9)
[…]
Proof.
$\mathrm{Var}[r]=\mathbb{E}\left[r-\mathbb{E}\left[r\right]\right]^{2}=\mathbb{E}\left[r^{2}\right]-\mathbb{E}^{2}\left[r\right]=\frac{\int^{M}_{K}r^{2}\exp(r/\alpha)dr}{\int^{M}_{K}\exp(r/\alpha)dr}-\left(\frac{\int^{M}_{K}r\exp(r/\alpha)}{\int^{M}_{K}\exp(r/\alpha)}\right)^{2}\;.$ (10)
Note that we can expand the integral at $\alpha=0$ with a Taylor series:
$\mathrm{Var}[r]=\frac{e^{(K+M)/\alpha}(-2\alpha^{2}+O(\alpha^{6}))+\left(e^{(2K)/\alpha}+e^{(2M/\alpha)}\right)\left(\alpha^{2}+O(\alpha^{6})\right)+(K-M)^{2}(-e^{(K+M)/\alpha})}{\left(e^{(K/\alpha)}-e^{(M/\alpha)}\right)^{2}}\;.$ (11)
Appendix C Datasets and Metrics
To evaluate MyGO’s effectiveness on complex question answering, we employ four representative datasets: 2WikiMultihopQA (2Wiki) ho2020constructing, QALD-10 QALD-10, HotpotQA yang2018hotpotqa, and MuSiQue trivedi2021musique. 2Wiki utilizes both structured (Wikidata) and unstructured (Wikipedia) data to create multi-hop questions. QALD-10 was initially designed for Multilingual Question Answering over KG but late was widely used as a benchmark on KBQA ToG. The other two datasets focus on multi-document and multi-hop reasoning. HotpotQA features complex questions requiring reasoning across multiple documents. MuSiQue consists of questions demanding two to four reasoning hops, constructed from five existing single-hop QA datasets.
Appendix D Implementation Details
D.1. Sampling Strategies
Reward Criterion
Note that some works utilize more complicated non-standard criteria, e.g., LLM-as-Judge ReSearch; sun2025zerosearch, Cover Exact Match (CEM) sun2025zerosearch; song2025r1. Since this is not the main focus of this work, we simply follow the standard criterion widely used in previous works.
Progressive Selection Threshold
Multiple-Attempt Sampling
For a batch of $n$ questions, we would iteratively sample trajectories for them. For a question, the sampling procedure will stop if $m$ satisfying trajectories are sampled or reach the predetermined maximum sampling attempts. With vLLM engine kwon2023efficient, we could accelerate the sampling procedures in parallel efficiently with multi-threading/multi-processing.
D.2. Training
Online Off-policy Optimization
In MyGO, we primarily focus on online off-policy optimization. Given a total of $N$ data examples (i.e., questions), we divide them into $T$ batches for $T$ update iterations. For each iteration, we sample trajectories (including answers) from $N/T$ questions to serve as pseudo-labels and optimize the models via MLE through a standard SFT pipeline (e.g., LLaMA-Factory zheng2024llamafactory). After completing one update iteration, we proceed to the next sampling iteration.
Trajectory Selection for Training.
In the Mujica framework, the planner and workers collaborate to reach the final answer. This interaction generates a single planner trajectory and multiple worker trajectories. It is important to clarify that in Section 3.3, the term ”trajectory” specifically refers to the planner’s trajectory. In contrast, the dialogue history for each worker is straightforward, comprising a single-turn interaction with the environment. The environment’s message to the worker consists of the question from the planner and passages from the retriever, while the worker’s message contains its response. If the reward associated with a planner trajectory exceeds a predefined threshold, that trajectory and its corresponding worker trajectories are selected as independent data for training.
Training Advances from Pre-training/SFT.
D.3. Hyperparameters and Configurations
Sampling.
Warmup.
Configuration.
Following the settings of regular SFT, we employ the AdamW optimizer loshchilov2019decoupled with a Cosine Annealing learning rate scheduler loshchilov2017sgdr. For the warmup phase, we configure the initial learning rate at 1e-5 and the minimum learning rate at 1e-6. In the MyGO training procedure, we set the initial learning rate to 5e-7 and the minimum learning rate to 1e-8. We leverage vLLM kwon2023efficient as our inference engine and LLaMA-Factory zheng2024llamafactory as our training framework.
Appendix E Further Study: Offline MyGO?
@@
Figure 4. F1 score on 2Wiki-KG training
Figure 5. F1 score on 2Wiki-Text training
Table 8. Mujica+MyGO w/ Qwen Online v.s. Offline on 2Wiki-KG and 2Wiki-Text Environments.
@@
We note that our goal is not to compare online and offline RL, as this has been discussed in many prior works tang2024understanding. Instead, we just aim to present an alternative approach for implementing MyGO.
@@
Appendix F Extended Related Works
@@
Retrieval-based methods dominate the landscape of textual MHQA since supporting contexts from external corpora can effectively compensate for deficiencies in the inherent knowledge of language models and mitigate hallucinations. Traditional approaches primarily utilized sentence transformers bgem3 for dense retrieval or BM25 BM25 for keyword matching. Recently, extensive research has focused on optimizing the QA pipeline, particularly through the decomposition of complex questions IM-RAG; ReAct and the iterative retrieval of relevant evidences in multiple steps SelfRAG; ReSP; Iter-RetRAG; IRCoT. Most notably, ReAct ReAct integrates retrieval actions into the reasoning trajectory exploration process, enabling dynamic interaction with external environments. IM-RAG IM-RAG optimizes multi-round retrieval and reasoning via reinforced learning. Search-o1 Search-o1 introduces an agentic RAG framework, empowering the language model to decide when to retrieve external knowledge and to refine retrieved contexts, achieving state-of-the-art performance.
@@
Early approaches lan-jiang-2020-query; SPARQL2021; KD-CoT formulates KBQA as a semantic parsing task, focusing on extracting entities and relations from natural language questions and translating them into SPARQL statements capable of executing logical queries over KGs TIARA; SymKGQA. While these methods provide accurate and interpretable answers, they encountered significant limitations since SPARQL statements generated by language models often suffer from syntactic errors, including incorrect entity and relation labels as well as unexpected grammatical mistakes DecAF; RoG; DoG. Moreover, the incomplete nature of KGs KGR3 can also lead to query failures, preventing these approaches from outputting correct answers.
@@
Recently, researchers have shifted towards treating KBQA as an information retrieval (IR) task, leveraging RAG and the advanced semantic understanding capabilities of LLMs to generate answers based on factual knowledge stored in KGs RRA. LLM-based KBQA methods can be generally categorized into two types: iterative exploration and subgraph reasoning. Iterative exploration methods ToG; ToG2; ToG-I; KG-CoT; PoG; FiDeLiS; GoG; KARPA; Paths-over-Graph; BeamRetrieval; DoG adopt a beam search-like strategy to retrieve the most relevant relations and neighboring entities, followed by pruning operations guided by LLMs. These methods exhibit strong traceability, enabling responsible and interpretable reasoning, but often come with the drawback of high computational complexity FastToG. Moreover, the absence of ground truth reasoning paths and intermediate nodes in most KBQA datasets limits the feasibility of applying SFT RoG; GNN-RAG. As a result, their performance is inevitably constrained by the inherent reasoning capabilities of LLMs. In contrast, subgraph reasoning methods RoG; GNN-RAG; SubgraphRAG; READS; GCR retrieve all candidate reasoning paths related to the question from the local neighborhood of the topic entities and guide the LLM to generate answers based on these paths in a single step. This approach enable direct improvement of the LLM’s in-context learning capabilities through SFT. Nevertheless, its effectiveness is often constrained by the quality of retrieved context, which typically cover only a subset of the knowledge required to address complex questions, omitting other critical information PoG. Furthermore, due to the incompleteness of KGs, the existence of relevant reasoning paths within the subgraph is not guaranteed. In such scenarios, LLM may hallucinate and produce incorrect answers.
Simplifying Reinforcement Learning for LLM postraining.
@@
Recent research has focused on simplifying RL algorithms for LLMs. For instance, GRPO shao2024deepseekmath simplifies PPO by replacing the value model with standard deviation. Similarly, ReMax li2024remax adapts VPO by using the logits of greedy decoding as a baseline for GAE. Other VPO modifications include RLOO ahmadian2024back, REINFORCE++hu2025reinforce++, and GPG chu2025gpg. Another line of research replaces VPO with MLE through advanced sampling techniques. Within this category, RAFT zhang2024raftadaptinglanguagemodel selects the highest-reward trajectory within a group, while RAFT++ xiong2025minimalist enhances this by incorporating rejection sampling. Both RAFT and RAFT++ can be viewed as specific instances of STaR zelikman2022star. MyGO aligns with this latter research direction. It selects trajectories using either a fixed or progressively improving threshold, a property ensuring that sampling approaches the optimal policy. We further provide a mathematical theory to explain the efficacy of this streamlined method. If we continue to simplify MyGO, it will converge to STaR zelikman2022star. However, our primary motivation stems from the goal of simplifying RL algorithms specifically for LLM agent training.
Appendix G Working Examples
@@
Table 9. Prompt template of Planner for text-based multi-hop question answering
@@
Before answering, outline your plan and reasoning process within ‘<think>’ tags. This includes understanding the question, identifying needed information, and planning your search queries.
@@
Please put your search query in ‘<search>’ tags, e.g., ‘<search>Who is the mother of Xawery Żuławski? </search>’
Every ‘<search>’ tag only accepts one search query.
But you can use multiple ‘<search>’ tags if needed.
@@
For example, if you directly search ‘Who is the mother of the director of film Polish-Russian War’, the search engine cannot find the information directly.
@@
Table 10. Prompt Template of Worker for Text-based Multi-hop Question Answering
@@
$[0]$ Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University. She made her acting debut in 1984, and has since become active in television dramas, most notably as the arrogant and ambitious Lady Choi in the 2003 period drama ”Dae Jang Geum” (or ”Jewel in the Palace”), which was a hit not only in Korea but throughout Asia.
$[1]$ Shin Kyeong-nim was born on April 6, 1936 in North Chungcheong Province, South Korea. When he was young, Shin Kyong-rim frequented the people of Koreaś rural villages and collected the traditional songs they sang. Much of his poetry represents a modernization of things he heard then Shin Kyeong-nim graduated in English Literature from Dongguk University, from which time he strove to become a creative writer. In 1955 and 1956, he made his formal literary debut with the publication of poems ”Day Moon” (Natdal), ”Reeds” (Galdae) and ”Statue of Stone” (Seoksang). He taught elementary school in his hometown for a period of time, before moving to Seoul to work as an editor for Hyundae munhak and Donghwa Publishers. But his strong desire to create poetry continued.
@@
$[3]$ Han Seung-yeon was born on July 24, 1988, in Seoul, South Korea. She made her acting debut as a child actress in a bit part in ”Dear Ends” (1993), ”Summer Showers” (1995) and ”Star in My Heart” (1997). She later left South Korea to study at Tenafly High School in New Jersey, United States. However, she withdrew from high school mid-course in order to pursue a singing career. After returning to South Korea, she debuted as a member of girl group Kara on March 29, 2007. During her time with the group, she passed a high school qualification exam, the College Scholastic Ability Test, and was accepted by Kyung Hee University, majoring in theater and film.
@@
Passage [0] clearly states that ”Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University.”
@@
Table 11. Prompt Template of Planner for Knowledge-based Question Answering
@@
Please think thoroughly and label the final answer with <answer>*** </answer>.
@@
Search([0], ”When did Blind Shaft come out?”)
Search([1], ”When did The Mask Of Fu Manchu come out?”)
@@
Table 12. Prompt Template of Worker for Knowledge-based Question Answering
@@
First, analyze the question to identify the key entity and requested information (e.g., ”mother,” ”birth date”).